Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability help

Expert replies
by jzw » Fri Feb 24, 2012 1:00 pm
So there these two questions.

"A six sided die with faces numbered one through six is rolled twice. What is the probability that the face with the number 2 on it will NOT be facing upwards on either roll?"

So I took 5/6 x 5/6 and got 25/36. That's the correct answer. Fine. And now the next question...

"Bob flips a coin four times. What is the probablity that he gets heads on at least one of the four flips?"

So why here are we supposed to take the probability that it does happen (1/2 x 1/2 x 1/2 x 1/2 = 1/16) and subtract that from 1 leaving 15/16? I get that the probability something will happen is 1 minus the probability that it won't happen. But then, if we do that for the first problem (with the six sided die) then we would take 1/6 x 1/6 = 1/36 and then if subtract that from 1 we are left with 35/36 which is NOT the correct answer.

Simply put - I am not sure when we are supposed to take the probability of the opposite thing happening and subtract that from one, and when we're not. PLEASE HELP!! - WAIT - is the answer because of the words "at least"?
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Fri Feb 24, 2012 2:26 pm
jzw wrote:So there these two questions.

"A six sided die with faces numbered one through six is rolled twice. What is the probability that the face with the number 2 on it will NOT be facing upwards on either roll?"

So I took 5/6 x 5/6 and got 25/36. That's the correct answer. Fine. And now the next question...

"Bob flips a coin four times. What is the probablity that he gets heads on at least one of the four flips?"

So why here are we supposed to take the probability that it does happen (1/2 x 1/2 x 1/2 x 1/2 = 1/16) and subtract that from 1 leaving 15/16? I get that the probability something will happen is 1 minus the probability that it won't happen. But then, if we do that for the first problem (with the six sided die) then we would take 1/6 x 1/6 = 1/36 and then if subtract that from 1 we are left with 35/36 which is NOT the correct answer.

Simply put - I am not sure when we are supposed to take the probability of the opposite thing happening and subtract that from one, and when we're not. PLEASE HELP!! - WAIT - is the answer because of the words "at least"?
P(good) + P(bad) = 1.

Given any probability question, we can determine P(good) or 1-P(bad).
The result is the same.
We should use whichever approach is easier and/or faster.

In 2 rolls, P(not 2 on either roll):
Good outcome: not 2 on the first roll and not 2 on the second roll.
Bad outcomes: 2 only on the first roll, 2 only on the second roll, 2 on both rolls.
Since there is only one good outcome but several bad outcomes, P(good) is faster.

In 4 coin flips, P(at least one heads):
Good outcomes: 1 heads, 2 heads, 3 heads, or 4 heads.
Bad outcome: all tails.
Since there is only one bad outcome, 1-P(bad) is faster.
But then, if we do that for the first problem (with the six sided die) then we would take 1/6 x 1/6 = 1/36 and then if subtract that from 1 we are left with 35/36 which is NOT the correct answer.
Your calculations here don't account for all of the bad outcomes listed above:
P(2 only on the first roll) = 1/6 * 5/6 = 5/36.
P(2 only on the second roll) = 5/6 * 1/6 = 5/36.
P(2 on both rolls) = 1/6 * 1/6 = 1/36.
Thus, P(not 2 on either roll) = 1 - (5/26+5/36+1/6) = 25/36.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by jzw » Fri Feb 24, 2012 5:02 pm
I'm not understanding how (conceptually) to figure it out the "longer" way in the six sided die question. Ie, if I was to change the question to "what are the chances the number 2 will land face up both times?" wouldn't the answer be 1/6 * 1/6 = 1/36 ?
Join the discussion

by GMATGuruNY » Fri Feb 24, 2012 5:20 pm
jzw wrote:I'm not understanding how (conceptually) to figure it out the "longer" way in the six sided die question. Ie, if I was to change the question to "what are the chances the number 2 will land face up both times?" wouldn't the answer be 1/6 * 1/6 = 1/36 ?
Step 1: What's a bad outcome?
NOT getting 2 both times.

Step 2: What are all of the ways to NOT get 2 BOTH times?
2 on the first roll but not on the second roll
2 on the second roll but not on the first roll
2 on neither roll.

Step 3: Are there more ways to get a good outcome or more ways to get a bad outcome?
In this case, there is only one way to get a good outcome (2 on each roll) but three ways to get a bad outcome, so it's quicker to determine P(good).

Here are other examples:

https://www.beatthegmat.com/combinations ... 02081.html

https://www.beatthegmat.com/baseball-t69721.html

https://www.beatthegmat.com/princeton-re ... 78922.html

https://www.beatthegmat.com/probability- ... 83693.html
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion